OpenStudy (anonymous):

Let y=2x^2+2x+2. Find the differential dy when x=1 and dx=0.1 Find the differential dy when x=1 and dx=0.2

OpenStudy (anonymous):

dy/dx=4x+2 multiply both side by dx dy=(4x+2)dx we know value of x and dx, so plug them in dy=(4(1)+2)(0.1) dy=.6